### v1

# Dirichlet dynamical zeta function for billiard flow

Abstract.
We study the Dirichlet dynamical zeta function $\eta_{D}(s)$ for billiard flow corresponding to several strictly convex disjoint obstacles. For large $\operatorname{Re}s$ we have $\eta_{D}(s)=\sum_{n=1}^{\infty}a_{n}e^{-\lambda_{n}s},\>a_{n}\in\mathbb{R}$ and $\eta_{D}$ admits a meromorphic continuation to $\mathbb{C}$. We obtain some conditions of the frequencies $\lambda_{n}$ and some sums of coefficients $a_{n}$ which imply that $\eta_{D}$ cannot be prolonged as entire function.
1. Introduction
Let $D_{1},\dots,D_{r}\subset\mathbb{R}^{d},\>{r\geqslant 3},\>d\geqslant 2,$ be compact strictly convex disjoint obstacles with $C^{\infty}$ smooth boundary and let $D=\bigcup_{j=1}^{r}D_{j}.$ We assume that every $D_{j}$ has non-empty interior and throughout this paper we suppose the following non-eclipse condition
$D_{k}\cap{\rm convex}\>{\rm hull}\>(D_{i}\cup D_{j})=\emptyset,$ (1.1)
for any $1\leqslant i,j,k\leqslant r$ such that $i\neq k$ and $j\neq k$. Under this condition all periodic trajectories for the billiard flow in $\Omega=\mathbb{R}^{d}\setminus\mathring{D}$ are ordinary reflecting ones without tangential intersections to the boundary of $D$. We consider the (non-grazing) billiard flow $\varphi_{t}$ (see [Pet24, Section 2] for the definition). Next the periodic trajectories will be called periodic rays. For any periodic ray $\gamma$, denote by $\tau(\gamma)>0$ its period, by $\tau^{\sharp}(\gamma)>0$ its primitive period, and by $m(\gamma)$ the number of reflections of $\gamma$ at the obstacles. Denote by $P_{\gamma}$ the associated linearized Poincaré map (see [PS17, Section 2.3] for the definition).
Let $\mathcal{P}$ be the set of all oriented periodic rays. The counting function of the lengths of primitive periodic rays $\Pi$ satisfies
$\sharp\{\gamma\in\Pi:\>\tau^{\sharp}(\gamma)\leqslant x\}\sim\frac{\mathrm{e}^{hx}}{hx},\quad x\to+\infty,$ (1.2)
for some $h>0$ (see for instance, [PP90, Theorem 6.5] for weak-mixing suspension symbolic flows). Thus there exists an infinite number of primitive periodic trajectories and for every small $\epsilon>0$ we have the estimate
$e^{(h-\epsilon)x}\leq\sharp\{\gamma\in\mathcal{P}:\>\tau(\gamma)\leqslant x\}\leqslant\mathrm{e}^{(h+\epsilon)x},\>x>C_{\epsilon}.$ (1.3)
Moreover, for some positive constants $C_{1},d_{1},d_{2}$ we have (see for instance [Pet99, Appendix])
$C_{1}\mathrm{e}^{d_{1}\tau(\gamma)}\leqslant|\det(\mathrm{Id}-P_{\gamma})|\leqslant\mathrm{e}^{d_{2}\tau(\gamma)},\quad\gamma\in\mathcal{P}.$ (1.4)
By using these estimates, we define for $\operatorname{Re}(s)\gg 1$ the Dirichlet dynamical zeta function $\eta_{D}(s)$ by
$\eta_{\mathrm{D}}(s)=\sum_{\gamma\in\mathcal{P}}(-1)^{m(\gamma)}\frac{\tau^{\sharp}(\gamma)\mathrm{e}^{-s\tau(\gamma)}}{|\det(\mathrm{Id}-P_{\gamma})|^{1/2}},$
where the sums run over all oriented periodic rays. This zeta function is important for the analysis of the distribution of the scattering resonances related to the Laplacian in $\mathbb{R}^{d}\setminus\bar{D}$ with Dirichlet boundary conditions on $\partial D$ (see [CP22, §1] for more details). Denote by $\sigma_{a}\in\mathbb{R},\sigma_{c}\in\mathbb{R}$ the abscissa of absolute convergence and the abscissa of convergence of $\eta_{D}$ respectively.
It was proved in [CP22, Theorem 1 and Theorem 4] that $\eta_{D}$ admits a meromorphic continuation to $\mathbb{C}$ with simple poles and integer residues. On the other hand, for $d=2$ [Sto01] and for $d\geqslant 3$ under some conditions [Sto12] Stoyanov proved that there exists $\varepsilon>0$ such that $\eta_{\mathrm{D}}(s)$ is analytic for $\operatorname{Re}s\geqslant\sigma_{a}-\varepsilon$.
There is a conjecture that $\eta_{D}$ cannot prolonged as entire function. This conjecture was established for obstacles with real analytic boundary (see [CP22, Theorem 3]) and for obstacles with sufficiently small diameters [Ika90b], [Sto09] and $C^{\infty}$ smooth boundary. If $\eta_{D}(s)$ is not an entire function, then we obtain two important corollaries:
(i) $\eta_{D}$ has infinite number of poles in some strip $\{z\in\mathbb{C}:\operatorname{Re}z\geq\beta\}$ (see [Pet24, Section 3] for a lower bound of the counting function of poles),
(ii) The modified Lax-Phillips conjecture (MLPC) for scattering resonances introduced by Ikawa [Ika90a] holds. (MLPC) says that there exists a strip $\{z\in\mathbb{C}:0<\operatorname{Im}z\leq\alpha\}$ containing an infinite number of scattering resonances for Dirichlet Laplacian in $\mathbb{R}^{d}\setminus\bar{D}$ (see [CP22, Section 1] for definitions and more precise results).
Let $\rho\in C_{0}^{\infty}(\mathbb{R};\mathbb{R}_{+})$ be an even function with ${\rm supp}\>\rho\subset\left[-1,1\right]$ such that $\rho(t)>1\quad\text{if}\quad|t| \leqslant 1/2.$ Let $(\ell_{j})_{j\in\mathbb{N}}$ and $(m_{j})_{j\in\mathbb{N}}$ be sequences of positive numbers such that $\ell_{j}\geqslant d_{0}=\min_{k\neq m}{\rm dist}\>(D_{k},D_{m})>0,\>m_{j}\geqslant\max\{1,\frac{1}{d_{0}}\}$ and let $\ell_{j}\to\infty,\>m_{j}\to\infty$ as $j\to\infty$. Set $\rho_{j}(t)=\rho(m_{j}(t-\ell_{j})),\quad t\in\mathbb{R},$ and introduce the distribution $\mathcal{F}_{\mathrm{D}}(t)\in{\mathcal{S}}^{\prime}(\mathbb{R}^{+})$ by
$\mathcal{F}_{\mathrm{D}}(t)=\sum_{\gamma\in\mathcal{P}}\frac{(-1)^{m(\gamma)}\tau^{\sharp}(\gamma)\delta(t-\tau(\gamma))}{|\det(I-P_{\gamma})|^{1/2}}.$
We have the following
Proposition 1.1.
The function $\eta_{\mathrm{D}}(s)$ cannot be prolonged as an entire function of $s$ if and only if there exists $\alpha_{0}>0$ such that for any $\beta>\alpha_{0}$ we can find sequences $(\ell_{j}),(m_{j})$ with $\ell_{j}\nearrow\infty$ as $j\to\infty$ such that for all $j\geqslant 0$ one has $e^{\beta\ell_{j}}\leqslant m_{j}\leqslant\mathrm{e}^{2\beta\ell_{j}}$ and
$|\langle\mathcal{F}_{\mathrm{D}},\rho_{j}\rangle|\geqslant\mathrm{e}^{-\alpha_{0}\ell_{j}}.$ (1.5)
More precisely, if $\eta_{D}$ cannot be prolonged as entire function, the existence of sequences $(\ell_{j}),\>(m_{j})$ with the above properties has been proved by Ikawa [Ika90a, Prop.2.3], while in the proof of Theorem 1.1 in [Pet24] it was established that if such sequences exist, the function $\eta_{D}$ has an infinite number of poles.
The conditions of Proposition 1.1 are difficult to verify. The purpose of this Note is to find another conditions which imply that $\eta_{D}$ cannot be prolonged as entire function. For this purpose we exploit the local trace formula (see [Pet24, Theorem 2.1]) and the summability by typical means of Dirichlet series introduced by Hardy and Riesz [HR64] (see also [DS22, Section 2]). It is convenient to write $\eta_{D}(s)$ as a Dirichlet series
$\eta_{D}(s)=\sum_{n=1}^{\infty}a_{n}e^{-\lambda_{n}s},\>\operatorname{Re}s\gg 1,$ (1.6)
where the frequencies are arranged as follows
$0<\lambda_{1}<\lambda_{2}<...<\lambda_{n}<...$
and
$a_{n}=\sum_{\gamma\in\mathcal{P},\tau(\gamma)=\lambda_{n}}\frac{(-1)^{m(\gamma)}\tau^{\sharp}(\gamma)}{|\det(\mathrm{Id}-P_{\gamma})|^{1/2}}.$ (1.7)
Our main result is the following
Theorem 1.1.
Suppose $\sigma_{c}<0$. Assume that there exist constants $C>0,\>\delta>h+1,\>\gamma<\sigma_{c}$ and an increasing sequence $m_{j}\nearrow\infty$ such that
$\lambda_{m_{j}}-\lambda_{m_{j}-1}\geq Ce^{-\delta\lambda_{m_{j}}},$ (1.8)
$|\sum_{n\geq m_{j}}a_{n}|\geq e^{\gamma\lambda_{m_{j}}}.$ (1.9)
Then $\eta_{D}(s)$ cannot be prolonged as entire function.
The condition $\sigma_{c}<0$ is not a restriction since if $\sigma_{c}\geq 0,$ the Dirichlet series
$\eta_{D}(s+\sigma_{c}+1)=\sum_{n}(a_{n}e^{-\lambda_{n}(\sigma_{c}+1)})e^{-\lambda_{n}s}=\sum_{n}b_{n}e^{-\lambda_{n}s}$
is convergent for $\operatorname{Re}s>-1$, hence it has a negative abscissa of convergence $\sigma_{b}$ and $\eta_{D}(s+\sigma_{c}+1)$ is entire if and only if $\eta_{D}(s)$ is entire. Moreover, in the proof of Theorem 1.1 (see section 3), assuming $\eta_{D}$ entire, one has the property
$\forall A<\sigma_{c},\>\exists C_{A}>0,\>\>|\eta_{D}(s)|\leq C_{A}(1+|\operatorname{Im}s|),\>\operatorname{Re}s\geq A$
which is satisfied also for $\eta_{D}(s+\sigma_{c}+1)$ with another constants $B_{A}$. Thus we may apply Theorem 1.1 if instead of (1.9) we have the condition
$|\sum_{n\geq m_{j}}b_{n}|\geq e^{\gamma_{1}\lambda_{m_{j}}},\>\gamma_{1}<\sigma_{b}.$ (1.10)
The conditions (1.8), (1.9) are satisfied if Bohr condition (see for instance, [DS22, §3.13])
$(BC)\>\>\>\>\>\>\exists C_{1}>0,\>\exists\ell>0,\>\>\forall n>0,\>\lambda_{n+1}-\lambda_{n}\geq C_{1}e^{-\ell\lambda_{n}}$
holds. Indeed, it is well known that in the case $\sigma_{c}<0$, one has the representation
$\sigma_{c}=\limsup_{n\to\infty}\frac{\log|\sum_{n\geq m}a_{n}|}{\lambda_{m}}.$
For small $\epsilon>0$ this implies the existence of a sequence $m_{j}\nearrow\infty$ such that
$|\sum_{n\geq m_{j}}a_{n}\geq e^{(\sigma_{c}-\epsilon)\lambda_{m_{j}}}$
and we obtain (1.9) with $\gamma=\sigma_{c}-\epsilon.$ The condition (BC) is very restrictive. The advantage of Theorem 1.1 is that (1.8) is always satisfied (see Section 3) for infinite number of frequencies $\lambda_{m_{j}-1},\>\lambda_{m_{j}}$ and the separation of some frequencies $\lambda_{m_{j}}$ only on the left is less restrictive than a separation of all frequencies on both sides. Applying Theorem 1.1, we obtain the following
Corollary 1.1.
Suppose $\sigma_{c}<0.$ Then if
$\liminf_{m\to\infty}\frac{\log|\sum_{n\geq m}a_{n}|}{\lambda_{m}}>-\infty,$ (1.11)
the function $\eta_{D}(s)$ cannot be prolonged as entire function.
In Section 4 for $\delta>h+2$ we study intervals $I(\lambda_{k},\delta)\subset[b,b+1],\>b\geq b_{0}$ with clustering frequencies and we obtain Corollary 4.1. We have infinite number of such intervals. Moreover, under some geometrical assumptions described in [PS12, Section 8] the number of such interval is exponentially increasing when $b\to\infty.$ Finally, assuming that the coefficients $a_{n}$ have lower bound (4.4) we show that for every interval $I(\lambda_{k},\delta)$ we have 4 possibilities concerning the behaviour of the corresponding sums. For 3 of these 4 possibilities it is possible to find frequencies satisfying (1.8), (1.9) (see Proposition 4.1).
The paper is organised as follows. In Section 2 we recall the local trace formula for $\eta_{D}.$ Assuming $\eta_{D}$ entire, we deduce the estimates (2.3). This makes possible to prove that that the abscissa of $k-$ summability $\sigma_{k}$ of $\eta_{D}$ is .$\infty$.- ∞ . In Section 3 we prove Theorem 1.1. Section 4 is devoted to intervals $I(\lambda_{k},\delta)$ with clustering frequencies and the constructions of frequencies satisfying (1.8) and (1.9).
2. Summation by typical means of $\eta_{D}$
In this section we apply the results of [DG16], [JT23, §6.1] and [Pet24] for vector bundles. For our exposition we need only the local trace formula containing the poles of the meromorphic continuation of cut off resolvents ${\bf 1}_{\tilde{V}_{u}}(-i\mathbf{P}_{k,\ell}-s)^{-1}{\bf 1}_{\tilde{V}_{u}}$ of some operators
[…]
Here $\tilde{V}_{u}$ us a neighborhood of the trapping set $\tilde{K}_{u}.$ The precise definitions of $\mathbf{P}_{k,\ell,q},\>\tilde{K}_{u}$ and the corresponding setting are complicated and they are not necessary for the analysis below. We refer to [Pet24, Section 2] for definitions and more details. Denote by ${\rm Res}\>(-i\mathbf{P}_{k,\ell,q})$ the set of the poles of the meromorphic continuation of the corresponding cut off resolvents.
For every $A>0$ and any $0<\epsilon\ll 1$ we have the following local trace formula (see [Pet24, Theorem 2.1])
$\sum_{k=0}^{d}\sum_{\ell=0}^{d^{2}-d}\sum_{\mu\in{\rm Res}\>(-i\mathbf{P}_{k,\ell,2}),\operatorname{Im}\mu>-A}(-1)^{k+\ell}e^{-i\mu t}$
$-\sum_{k=0}^{d}\sum_{\ell=0}^{d^{2}-d}\sum_{\mu\in{\rm Res}\>(-i\mathbf{P}_{k,\ell,1}),\operatorname{Im}\mu>-A}(-1)^{k+\ell}e^{-i\mu t}$ (2.1)
$+F_{A}(t)=\mathcal{F}_{D}(t),\>t>0.$
Here $F_{A}(t)\in{\mathcal{S}}^{\prime}(\mathbb{R})$ is supported in $(0,\infty)$, the Laplace-Fourier transform $\hat{F}_{A}(\lambda)$ of $F_{A}(t)$ is holomorphic for $\operatorname{Im}\lambda<A-\epsilon$ and satisfies the estimate
$|\hat{F}_{A}(\lambda)|={\mathcal{O}}_{A,\epsilon}(1+|\lambda|)^{2d^{2}+2d-1+\epsilon},\>\operatorname{Im}\lambda<A-\epsilon.$ (2.2)
Notice that the poles in ${\rm Res}\>(-i\mathbf{P}_{k,\ell,q})$ are simples with positive integer residues [CP22]. For the terms with fixed $q$ the cancellations in (2) could appear only between the terms with $k+\ell$ odd and $k+\ell$ even. On the other hand, taking the difference of terms with $q=2$ and $q=1$ we obtain more cancelations.
If the following we assume that $\eta_{D}$ can be prolonged as entire function. Then
$\eta_{D}(-i\lambda)=\langle\mathcal{F}_{D},e^{it\lambda}\rangle=\sum_{\gamma\in\mathcal{P}}\frac{(-1)^{m(\gamma)}\tau^{\sharp}(\gamma)e^{i\lambda\tau(\gamma)}}{|\det(\operatorname{Id}-P_{\gamma})|^{1/2}},\>\operatorname{Im}\lambda\geq C\gg 1$
admits an analytic continuation for $\operatorname{Im}\lambda<C.$ For fixed $A>0$ the function $\eta_{D}(-i\lambda)$ has no poles $\mu$ with $\operatorname{Im}\mu>-A$ and in (2) all terms involving poles will be canceled. Consequently, from (2) we obtain
$\eta_{D}(-i\lambda)=\hat{F}_{A}(-\lambda),\>\operatorname{Im}\lambda>-A+\epsilon.$
Setting $-i\lambda=s=\sigma+it,\>\sigma\in\mathbb{R},\>t\in\mathbb{R}$, this implies
$|\eta_{D}(s)|\leq C_{A}(1+|s|)^{2d^{2}+2d-1}\leq B_{A}(1+|t|)^{2d^{2}+2d-1},\>\sigma\geq-A+\epsilon.$ (2.3)
Here we used the fact that $|\eta_{D}(s)|$ is bounded for $\sigma\geq C_{0}>0$ with sufficiently large $C_{0}>0$ and $|s|\leq\max\{A,C_{0}\}+|t|$ for $-A\leq\sigma\leq C_{0}.$ We may apply the above argument for every $A>0,$ so the bound (2.3) holds for every $A>0$ with constants $B_{A}$ depending of $A$. The crucial point is that the power $12d^{2}+2d-$ 1 is independent of $A$.
Applying the Phragmént- Lindelöf principle for entire function $\eta_{D}(s)$ in the strip
$\{z\in\mathbb{C}:-A+\epsilon\leq\operatorname{Re}z\leq C_{0}\},$
one deduces
$|\eta_{D}(\sigma+it)|\leq D_{\sigma,A}(1+|t|)^{\kappa(\sigma)},\>-A+\epsilon\leq\sigma\leq C_{0}$
with
$\kappa(\sigma)=\frac{C_{0}-\sigma}{C_{0}+A-\epsilon}(2d^{2}+2d-1),\>-A+\epsilon\leq\sigma\leq C_{0}.$
For fixed $\sigma$, taking $A$ sufficiently large we obtain for every small $0<\nu\ll 1$ the estimate
$|\eta_{D}(\sigma+it)|\leq B_{\sigma,\nu}(1+|t|)^{\nu},\>\sigma\leq C_{0}.$ (2.4)
Next we recall the summation by typical means of Dirichlet series (see [HR64, Section IV, §2], [DS22, Section 2] for more details). For $k>0$ consider
$C_{\lambda}^{k}(u)=\sum_{\lambda_{n}<u}(u-\lambda_{n})^{k}a_{n}e^{-\lambda_{n}s}.$
We say that the series $\sum_{n=1}^{\infty}a_{n}e^{-\lambda_{n}s}$ is $(\lambda,k)$ summable if
$\lim_{u\to\infty}\frac{C_{\lambda}^{k}(u)}{u^{k}}=C.$
There exists a number $\sigma_{k}$ such that the series is $(\lambda,k)$ summable for $\sigma>\sigma_{k}$ and not $(\lambda,k)$ summable for $\sigma<\sigma_{k}$ The number $\sigma_{k}$ is called abscissa of $k$- summability of the series. We have the following
Theorem 2.1 (Th.41, [HR64]).
Suppose that the series $f(s)=\sum_{n=1}^{\infty}a_{n}e^{-\lambda_{n}s}$ admits an analytic continuation for $\sigma>\eta.$ Suppose further that $k$ and $k^{\prime}$ are positive numbers such that $k^{\prime}<k$ and for all small $\delta$ we have
$|f(s)|\leq\mathcal{C}_{\delta}(1+|s|)^{k^{\prime}}$
uniformly for $\sigma\geq\eta+\delta>\eta.$ Then $f(s)$ is $(\lambda,k)$ summable for $\sigma>\eta.$
In fact, the above theorem in [HR64] is given without proof. The reader may consult Corollary 3.8 and Corollary 3.9 in [DS22] for a proof and other results related to Theorem 2.1 and $(\lambda,k)$ summability. The estimates (2.4) combined with Theorem 2.1 imply the following
Proposition 2.1.
[…]
3. Proof of Theorem $1.1$
Throughout this section we assume that $\eta_{D}(s)$ can be prolonged as entire function. Choose $\delta>h+2$. First, it is easy to see that in every interval $[b,b+1],b\geq b_{0}\gg 1$ we have subintervals $[\alpha,\beta]\subset[b+b+1]$ of length greater than $e^{-\delta b}$ which does not contain frequencies. It is sufficient to write $[b,b+1]$ as union of $e^{\delta b}$ intervals of length $e^{-\delta b}$ and to use the bounds (1.3).
We have the following simple
Lemma 3.1.
Fix a small $20<\epsilon<1/$ 2 . There exists $b_{0}\geq\max\{3/h,1\}$ depending of $\epsilon$ so that for $\alpha\geq b_{0}$ we have
$\sharp\{\gamma\in\Pi:\alpha\leq\tau^{\sharp}(\gamma)\leq\alpha+\epsilon\}>\frac{\epsilon(1-\eta)e^{\alpha h}}{3(\alpha+\epsilon)}.$ (3.1)
Proof.
Choose $0<\eta<\epsilon$ so that $4\eta\leq\frac{\epsilon h}{3(1+\epsilon)}.$ For $x\geq b_{0}(\eta)\gg 1$ the asymptotics (1.2), imply the estimates
$\frac{e^{hx}}{hx}(1-\eta)\leq\sharp\{\gamma\in\Pi:\tau^{\sharp}(\gamma)\leq x\}\leq\frac{e^{hx}}{hx}(1+\eta).$
Therefore for $\alpha\geq b_{0}(\eta)$ we obtain
$\sharp\{\gamma\in\Pi:\alpha\leq\tau^{\sharp}(\gamma)\leq\alpha+\epsilon\}\geq\frac{e^{h(\alpha+\epsilon)}}{h(\alpha+\epsilon)}(1-\eta)-\frac{e^{h\alpha}}{h\alpha}(1+\eta)$
$>\frac{(1-\eta)e^{\alpha h}}{h(\alpha+\epsilon)}\Bigl{[}1+\epsilon h-\frac{(\alpha+\epsilon)(1+\eta)}{\alpha(1-\eta)}\Bigr{]}.$
On the other hand, we have $\frac{1}{\alpha}\leq\frac{h}{3}$ and
$4\eta\leq\frac{\epsilon h}{3(1+\epsilon)}\leq\frac{\epsilon h}{3(1+\frac{\epsilon}{\alpha})}.$
Then
$\frac{(\alpha+\epsilon)(1+\eta)}{\alpha(1-\eta)}=\Bigl{(}1+\frac{\epsilon}{\alpha}\Bigr{)}\Bigl{(}1+\frac{2\eta}{1-\eta}\Bigr{)}\leq(1+\frac{\epsilon}{\alpha})(1+4\eta)\leq 1+\frac{2\epsilon h}{3}$
and one deduces (3.1). ∎
Proof of Theorem 1.1. We start with the formula of the abscissa of $k-$summability $\sigma_{k}<0$ in the case when $k\in\mathbb{N}$ is integer established by Kuniyeda [Kun16, Theorem E]. More precisely, we have
$\sigma_{k}=\limsup_{u\to\infty}\frac{\log|R^{k}(u)|}{u^{k}},$ (3.2)
where $R^{k}(u)=\sum_{\lambda_{n}>u}a_{n}(\lambda_{n}-u)^{k}.$ We are interesting of the case $k=1.$ Let $\delta,\gamma$ be the constants given in (1.8), (1.9), respectively. By Proposition 2.1 for $\eta_{D}(s)$ we have $\sigma_{1}=-\infty$. We fix $\gamma_{1}>0$ so that $-\gamma_{1}<-\delta+\gamma-1.$ Then (3.2) implies that there exists $M=M(\gamma_{1})>1$ such that
$|R(u)|=|\sum_{\lambda_{n}>u}a_{n}(\lambda_{n}-u)|\leq e^{-\gamma_{1}u},\>\forall u\geq M.$
Let
$\lambda_{m_{j}}-\lambda_{m_{j}-1}\geq Ce^{-\delta\lambda_{m_{j}}},\>\lambda_{m_{j}-2}\geq M,\>|\sum_{n\geq m_{j}}a_{n}|\geq e^{\gamma\lambda_{m_{j}}}.$
Moreover, for $M$ large by using (3.1), we have $\lambda_{m_{j}}-\lambda_{m_{j}-1}<1.$
Choose $u_{m_{j}-1},u_{m_{j}}$ so that $\lambda_{m_{j}-2}<u_{m_{j}-1}<\lambda_{m_{j}-1}<u_{m_{j}}<\lambda_{m_{j}}$ and write
$R(u_{m_{j}-1})-R(u_{m_{j}})=a_{m_{j}-1}(\lambda_{m_{j}-1}-u_{m_{j}-1})+(u_{m_{j}}-u_{m_{j}-1})\sum_{\lambda_{n}>u_{m_{j}}}a_{n}.$
We choose $\lambda_{m_{j}-1}-u_{m_{j}-1}=\epsilon_{j}\ll 1$ sufficiently small to arrange
$|a_{m_{j}-1}|(\lambda_{m_{j}-1}-u_{m_{j}-1})\leq e^{-\gamma_{1}\lambda_{m_{j}}}.$
(Exploiting (1.4), we obtain an upper bound $|a_{n}|\leq e^{c\lambda_{n}},\>\forall n$ with $c>0$ independent of $\lambda_{n}$, but this is not necessary for the estimation above.) Next
$u_{m_{j}}-u_{m_{j}-1}=(u_{m_{j}}-\lambda_{m_{j}})+(\lambda_{m_{j}}-\lambda_{m_{j}-1})+(\lambda_{m_{j}-1}-u_{m_{j}-1}).$
Taking $\epsilon_{j}$ very close to 0, if it is necessary, we choose $u_{m_{j}}=\lambda_{m_{j}}-\epsilon_{j},$ and deduce
$u_{m_{j}}-u_{m_{j}-1}=\lambda_{m_{j}}-\lambda_{m_{j}-1}\geq Ce^{-\delta\lambda_{m_{j}}}.$
Then
$Ce^{(-\delta+\gamma)\lambda_{m_{j}}}\leq(u_{m_{j}}-u_{m_{j}-1})|\sum_{n\geq m_{j}}a_{n}|$
$=|R(u_{m_{j}-1})-R(u_{m_{j}})-a_{m_{j}-1}(\lambda_{m_{j}-1}-u_{m_{j}-1})|$
$\leq e^{-\gamma_{1}u_{m_{j}-1}}+e^{-\gamma_{1}u_{m_{j}}}+e^{-\gamma_{1}\lambda_{m_{j}}}$
$\leq\Bigl{(}2e^{\gamma_{1}(\lambda_{m_{j}}-u_{m_{j}-1})}+1\Bigr{)}e^{-\gamma_{1}\lambda_{m_{j}}}.$
Since
$\lambda_{m_{j}}-u_{m_{j}-1}=\lambda_{m_{j}}-\lambda_{m_{j}-1}+\epsilon_{j}<3/2,$
the above inequality yields
$1\leq\frac{1}{C}\Bigl{(}2e^{\frac{3}{2}\gamma_{1}}+1\Bigr{)}e^{(-\gamma_{1}+\delta-\gamma)\lambda_{m_{j}}}$
and we obtain a contradiction for $\lambda_{m_{j}}\to\infty.$ This completes the proof.
Since (1.8) is always satisfied for suitable frequencies $\lambda_{m_{j}-1}\nearrow\infty,\lambda_{m_{j}}$ (see Section 4), exploiting (1.11) we may arrange the condition (1.9) for $\lambda_{m_{j}}$ large enough. An application of Theorem 1.1 yields Corollary 1.1.
4. Intervals with clustering frequencies
We fix $\delta>h+2$ and $e^{-b}<\epsilon\ll 1/2$ and consider an interval $[b,b+1],\>b\geq b_{0}(\epsilon).$ Let $\lambda_{k}\in[b+e^{-b},b+1-e^{-b}].$ To examine the clustering around $\lambda_{k}$, we construct some sets. Introduce
$J_{\delta}(\mu)=(\mu,\mu+\mathrm{e}^{-\delta b}).$
If $\lambda_{k+1}\notin J_{\delta}(\lambda_{k}),$ we stop the construction on the right. If $\lambda_{k+1}\in J_{\delta}(\lambda_{k})$, one considers $J_{\delta}(\lambda_{k+1}).$ In the case $\lambda_{k+2}\notin J_{\delta}(\lambda_{k+1})$, we stop the construction. Otherwise, we continue with $J_{\delta}(\lambda_{k+2})$ up to the situation when $\lambda_{k+q+1}\notin J_{\delta}(\lambda_{k+q}).$ It is clear that such $q$ exists.We repeat the same construction moving on the left introducing
$G_{\delta}(\mu)=(\mu-\mathrm{e}^{-\delta b},\mu).$
We stop when $\lambda_{k-p-1}\notin G_{\delta}(\lambda_{k-p}).$ Set $I(\lambda_{k},\delta)=[\lambda_{k-p},\lambda_{k+q}].$ The integers $p,q$ depend on $\lambda_{k}$, but we omit this in the notations below. Clearly, if we take another frequency $\lambda_{k^{\prime}}\in I(\lambda_{k},\delta),$ we obtain by the above construction the same interval. The number of the frequencies in $I(\lambda_{k},\delta)$ is bounded by $e^{(h+\epsilon)(b+1)}$ and
$\lambda_{k+q}-\lambda_{k-p}\leq e^{(h-\delta+\epsilon)b+(h+\epsilon)}<e^{-b},\>b\geq b_{0}(\epsilon).$ (4.1)
This estimate implies $\lambda_{k+q}<b+1,\>\lambda_{k-p}>b,$ so $I(\lambda_{k},\delta)\subset(b,b+1).$ By Lemma 3.1, the intervals without frequencies have lengths less than $\epsilon$. Let $M(\epsilon,\delta,b)$ be the number of the sets
$I(\lambda_{k},\delta)\cup(\lambda_{k+q},\lambda_{k+q+1}),\>e^{-\delta b}\leq\lambda_{k+q+1}-\lambda_{k+q}<\epsilon,\>\lambda_{k}\in[b+e^{-b},b+1-e^{-b}].$
Taking the union of such sets, we obtain
$M(\epsilon,\delta,b)(\epsilon+e^{-b})\geq 1-2\epsilon-2e^{-b}.$
Thus for large $b$ we have
$M(\epsilon,\delta,b)>\frac{1-2\epsilon-2e^{-b}}{\epsilon+e^{-b}}=\frac{1}{\epsilon}-2+\mathcal{O}_{\epsilon}(e^{-b}).$ (4.2)
Hence we have at least $\Bigl{[}\frac{1}{\epsilon}\Bigr{]}-2$ frequencies $\lambda_{m_{j}}\in[b+e^{-b},b+1-e^{-b}]$ with
$\lambda_{m_{j}-p_{j}}-\lambda_{m_{j}-p_{j}-1}>e^{-\delta\lambda_{m_{j}-p_{j}-1}},\>\lambda_{m_{j}+q_{j}+1}-\lambda_{m_{j}+q_{j}}>e^{-\delta\lambda_{m_{j}+q_{j}}},$ (4.3)
where $[a]$ denotes the entire part of $a$.
Now let $\gamma\gg 1$ be fixed. Given an interval $I(\lambda_{k},\delta)\subset(b,b+1),$ we have 2 possibilities:
$(i)\>\>\>|\sum_{n\geq k-p}a_{n}|\geq e^{-\gamma\lambda_{k-p}},\>\>(ii)\>\>\>|\sum_{n\geq k-p}a_{n}|<e^{-\gamma\lambda_{k-p}}.$
In the case (i) the conditions (1.8), (1.9) are satisfied for $\lambda_{k-p-1}$ and $\lambda_{k-p}.$ If one has (ii), and $|\sum_{n\geq k+q+1}a_{n}|<e^{-\gamma\lambda_{k+q+1}},$ by triangle inequality one deduces
$|\sum_{n=k-p}^{k+q}a_{n}|\leq e^{-\gamma\lambda_{k-p}}+e^{-\gamma\lambda_{k+q+1}}<2e^{-\gamma\lambda_{k-p}}.$
Thus if (ii) holds, and $|\sum_{n=k-p}^{k+q}a_{n}|\geq 2e^{-\gamma\lambda_{k-p}}$ the the conditions (1.8), (1.9) are satisfied for $\lambda_{k+q}$ and $\lambda_{k+q+1}.$ Taking into account (4.3) and applying Theorem 1.1, we obtain the following
Corollary 4.1.
Suppose $\sigma_{c}<0$. Suppose that there exist constants $\delta>h+2,\gamma\gg 1$ and a sequence of intervals
$I(\lambda_{m_{j}},\delta)=[\lambda_{m_{j}-p_{j}},\lambda_{m_{j}+q_{j}}],\>\lambda_{m_{j}}\nearrow\infty$
[…]
$|\sum_{n=m_{j}-p_{j}}^{m_{j}+q_{j}}a_{n}|\geq 2e^{-\gamma\lambda_{m_{j}-p_{j}}}.$
Then $\eta_{D}$ cannot be prolonged as entire function.
It is important to increase the number of intervals included in $[b,b+1]$ satisfying the property (4.3). By using Lemma 3.1, we see that for $\epsilon\searrow 0$ we have $b_{0}(\epsilon)\nearrow\infty$ so a more precise asymptotics for the counting functions of the number of frequencies with remainder is necessary. Under some geometrical assumptions, it was proved (see [PS12, Theorem 4]) that we may replace $\epsilon$ by $e^{-\mu b}$ with small $0<\mu<h$ and obtain a lower bound of
$\sharp\{\gamma\in\Pi:\alpha\leq\tau^{\sharp}(\gamma)\leq\alpha+e^{-\mu b}\}.$
These assumptions are satisfied for $d=2$, while for $d\geq 3$ one make some restrictions. We refer to [PS12, Section 8] for precise results and more details. Under these assumptions we will obtain $M(e^{-\mu b},\delta,b)\sim e^{\mu b}$ as $b\to\infty$ so one obtains exponentially increasing number of intervals satisfying (4.3). This increases the possibilities to satisfy the conditions of Theorem 1.1.
To obtain a lower bound for $|a_{n}|,\>\forall n\geq n_{0}$, introduce the condition
(L) There exist constants $c_{1}>0,c_{2}>0,$ independent of $n$ such that
$|a_{n}|\geq c_{1}e^{-c_{2}\lambda_{n}},\>\forall n\geq n_{0}.$ (4.4)
The condition (4.4) holds if the lengths of primitive periodic rays $\gamma\in\Pi$ are rationally independent, because (1.7) will contain only one term and from (1.4) we obtain (4.4) with $c_{1}=\min_{i\neq j}{\rm dist}\>(D_{i},D_{j})$ and $c_{2}=d_{2}/2.$ This rationally independence has been proved for generic domains (see [PS17, Theorem 6.2.3]). Then if (L) holds and
$|\sum_{k\geq m}a_{k}|<e^{-\gamma\lambda_{m}},\>|\sum_{k\geq m+1}a_{k}|<e^{-\gamma\lambda_{m+1}}$ (4.5)
with $\gamma>c_{2}+1$, one deduces $c_{1}e^{-c_{2}\lambda_{m}}\leq|a_{m}|<2e^{-\gamma\lambda_{m}}$ which is impossible for large $\lambda_{m}.$ Hence at least one of the estimates (4.5) does not hold.
Going back to interval $I(\lambda_{k},\delta)$, notice that for $\lambda_{k+q}$ one has also 2 possibilities:
$(iii)\>\>\>|\sum_{n\geq k+q}a_{n}|\geq e^{-\gamma\lambda_{k+q}},\>\>(iv)\>\>\>|\sum_{n\geq k+q}a_{n}|<e^{-\gamma\lambda_{k+q}}.$
Assuming (L) and $\gamma>c_{2}+1$, in the case (iv) the conditions (1.8), (1.9) are satisfied for $\lambda_{k+q}$ and $\lambda_{k+q+1}.$ Consequently, we obtain the following
Proposition 4.1.
[…]
A more fine analysis of the lower bound of the sums $\sum_{n=k_{j}-p_{j}}^{k_{j}+q_{j}}a_{n}$ should imply more precise results.